### v1

# Classification of three-qubit genuine entangled states using concurrence fill

Abstract
Despite the successful experimental generation and verification of genuine multipartite entanglement, several existing entanglement measures remain insufficient to reliably capture its presence. In this study, we overcome this challenge by utilizing a geometric measure known as concurrence fill, which quantifies genuine multipartite entanglement using the area associated with the underlying concurrence triangle. Firstly, the concurrence fill for three-qubit pure states is reformulated using three tangle and partial tangles. This yields a representation that is more tractable and operational. Using this formulation, we derive a criterion to classify GHZ and W class of states. Next, we analyse the rank-2 mixture of generalized GHZ and W class of states. Furthermore, we derive the concurrence fill for the eigenstates of the corresponding mixture and obtain an upper bound for mixed states with zero tangle.
Keywords: Quantum Entanglement. Genuine entangled states. Concurrence fill
I Introduction
Quantum entanglement is an important physical resource in quantum information theory, which distinguishes between quantum and classical worlds. Entanglement has applications in bipartite as well as multipartite quantum systems. Genuine multipartite entanglement (GME) is a significant concept for multipartite systems, since it plays an essential role in quantum communication karlson1998 ; yeo2006 and quantum cryptography hillery1999 . Recognizing the significance of entanglement as a valuable resource, it becomes natural to explore its quantification through theoretical methods. For the bipartite system, various entanglement measures have been developed, such as concurrence wootters1998 ; wootters1997 , negativity vidal2002 , entanglement of formation wootters1998 , etc. Carvalho et al. derived the generalization of concurrence, namely the generalized multipartite concurrence carvalho2004 . A novel entanglement monotone for genuine multipartite entanglement is proposed, serving as a natural extension of the negativity and coinciding with it in the bipartite case jungnitch2011 . Additionally, other bipartite entanglement measures, including the relative entropy of entanglement and the robustness of entanglement, are shown to be readily adaptable to the multipartite framework vidal2001 .
Acin et al introduced the classification of mixed three-qubit states into different entanglement classes acin2001 . These states are classified into separable, biseparable, W class and GHZ class of states. Any nontrivial tripartite entangled state can be converted, by means of SLOCC, into one of two standard forms, namely, GHZ or W state, which are given as follows dur2000 :
$|GHZ\rangle$ $=$ $\frac{1}{\sqrt{2}}\left(|000\rangle+|111\rangle\right)$ (1)
$|W\rangle$ $=$ $\frac{1}{\sqrt{3}}\left(|100\rangle+|010\rangle+|001\rangle\right)$ (2)
GHZ and W class of states are stochastically inequivalent classes of entangled states, i.e., a state in GHZ class cannot be converted into W class by stochastic LOCC.
A tripartite measure of entanglement known as three-tangle (or tangle) was introduced by Coffman, Kundu and Wootters (CKW) as residual entanglement ckw2000 . Tangle was derived as consequence of the most significant property of the quantum entanglement, i.e., its monogamous behavior. It plays a significant role in the quantum cryptography which restricts the information gain by an eavesdropper qaisar2017 , and in quantum channel discrimination sen2016 . Monogamy relations were first established by Coffman, Kundu and Wootters (CKW) in ckw2000 by considering a three-qubit system. It was shown that the entanglement between the subsystems $A$ and $B$ and entanglement between the subsystems $A$ and $C$ cannot be greater than the entanglement between the subsystems $A$ and $BC$. The concept of monogamy property has been extended to the general $N$-qubit system verstrate2006 . Monogamy inequalities for various entanglement measures such as concurrence ckw2000 , entanglement of formation olive2014 , negativity fan2007 ; kim2009 , Tsallis-q entanglement kim2010 ; luo2016 and Renyi-$\alpha$ entanglement sanders2010 have been further studied.
Quantification of entanglement for multipartite systems is a great challenge due to the much richer mathematical structures involved as compared to the bipartite case. A valid measure of multipartite entanglement must meet two key criteria to be considered a genuine multipartite entanglement (GME) measure. These criteria, identified by Ma et al. ma2011 , are: (a) the measure must be zero for all product and biseparable states, and (b) the measure must be positive for all nonbiseparable states (GHZ and W classes for the three-qubit case). In eberly2021 , authors used the monogamy inequality on entanglement distribution in three-qubit systems to introduce a new measure of genuine tripartite entanglement called concurrence fill. It is a GME measure since it is non-zero for GHZ as well as W class of states. The concurrence fill was studied in the experiments of three-flavor neutrino oscillations li2022 . The extension of concurrence fill to four partite systems, is presented in mishra , which represents concurrence fill as a combination of areas of cyclic quadrilateral and triangle structures, arising from two distinct types of bipartitions.
In our work, we study the concurrence fill for pure and mixed three-qubit states. Firstly, we provide an analytic reformulation of concurrence fill in terms of three-tangle and partial tangle. This formulation is constructive because it has been shown in literature that tangle as well as partial tangles can be measured experimentally datta2018 . Since tangle is zero for W class of states, it cannot classify W class and GHZ class of states. To overcome this problem we use our formulation of concurrence fill to devise a criterion that classifies GHZ and W class of states in three-qubit systems. In eberly2021 , authors have introduced the concurrence fill for three-qubit pure states. A corresponding measure for mixed states can be obtained using convex roof extension. This method involves minimizing the concurrence fill of a mixed state $\rho$ over all possible pure state decompositions of $\rho$. Hence, convex roof extension is a powerful but computationally challenging tool. In this work, we consider the family of rank-2 mixture of generalized GHZ and W states in three qubit system. Since all the elements in a decomposition are superpositions of the eigenstates of the given density matrix, the minimum value of concurrence fill will depend on all the possible concurrence fill values for those superpositions.We derive the concurrence fill for the eigenstates of the density matrices belonging to this family of rank-2 mixed states. Next, we deduce a lower bound of the concurrence fill of these eigenstates. Moreover, we derive an upper bound for the concurrence fill of mixed states in this family of states.
In section II, few entanglement measures such as concurrence, tangle and concurrence fill have been discussed. Section III contains the reformulation of concurrence fill in terms of tangle and partial tangle followed by criteria for classification of GHZ and W class of states. In section IV, we study the concurrence fill for rank-2 mixed states. Finally, we conclude in section V.
II Entanglement Measures
Several entanglement measures have been proposed, particularly for the bipartite states. Yet, the situation is much more complicated for multipartite systems. If an N-partite state is neither separable nor biseparable, it is genuine $N$-partite entangled. Any genuine multipartite entanglement measure (GME), denoted by $E(\rho)$ for the state $\rho$ satisfies the following necessary conditions vedral .
  • (i)
    $E(\rho)=0$ if and only if $\rho$ is separable or biseparable.
  • (ii)
    $E(\rho)>0$ for all genuinely entangled states.
  • (iii)
    $E(\rho)$ is independent under local unitary transformations
    $E(\rho)=U\rho U^{\dagger}$ (3)
    where $U$ is an arbitrary local unitary operator.
  • (iv)
    $E(\rho)\geq\sum_{i}p_{i}E(\rho_{i})$ where $\{p_{i},\rho_{i}\}$ represents the ensemble of states obtained by applying LOCC operations on $\rho$.
Firstly, let us investigate some entanglement measures that are closely related to concurrence wootters1998 ; wootters1997 ; ckw2000 .
II.1 Concurrence
Concurrence is a well-defined measure of entanglement. Let $\rho_{AB}$ be a density matrix describing a two-qubit pure or mixed state. Then the spin flipped density matrix $\tilde{\rho}_{AB}$ is defined as
$\tilde{\rho}_{AB}=(\sigma_{y}\otimes\sigma_{y})\rho_{AB}^{*}(\sigma_{y}\otimes\sigma_{y})$ (4)
where * denotes complex conjugation in the standard basis $\{|00\rangle,|01\rangle,|10\rangle,|11\rangle\}$; and $\sigma_{y}$ is the Pauli matrix given as
$\begin{pmatrix}0&-i\\ i&0\end{pmatrix}$ (5)
The concurrence of the state $\rho_{AB}$ is defined as
$C_{AB}=max\{\lambda_{1}-\lambda_{2}-\lambda_{3}-\lambda_{4},0\}$ (6)
where $\lambda_{i}^{\prime}s$ are the real eigenvalues of the positive semi-definite matrix ${\rho}_{AB}\tilde{\rho}_{AB}$. If the two qubit state $\rho_{AB}$ is a pure state, concurrence is defined as
$C_{AB}=2\sqrt{det\rho_{A}}$ (7)
where $\rho_{A}$ is the reduced state (trace of $\rho_{AB}$ over the qubit $B$).
II.2 Tangle
Tangle, or three-tangle, extends the idea of concurrence to tripartite systems. For a tripartite pure state $|\psi_{ABC}\rangle$, tangle is expressed as
$\tau(|\psi_{ABC}\rangle)=C^{2}_{A|BC}-C^{2}_{AB}-C^{2}_{AC}$ (8)
where $C^{2}_{A|BC}$ represents the amount of bipartite entanglement between qubit $A$ and the composite qubits $BC$ quantified by the Wootters’ concurrence wootters1998 . This is to note that tangle is unchanged by permutations of indices that enter in the considered bipartition. For GHZ states, tangle is maximal, i.e., $\tau(|GHZ\rangle)=1$ and it vanishes for all separable, biseparable and W-class of states. Also, tangle does not increase on average under LOCC.
The three-tangle has also been proposed for mixed states via convex-roof extension of pure states. For a mixed three-qubit state $\rho_{ABC}=\sum_{i}p_{i}|\psi_{i}\rangle\langle\psi_{i}|$, tangle is given as
$\tau(\rho_{ABC})=min_{\{p_{i},|\psi_{i}\rangle\}}\sum_{i}{p_{i}\tau(|\psi_{i}\rangle})$ (9)
where the minimum is taken over all convex decompositions of $\rho_{ABC}$.
II.3 Partial Tangle
To further characterize entanglement in three-qubit systems, a quantity known as partial tangle was introduced in slee2005 . It is defined as follows.
$\tau_{ij}=\sqrt{C^{2}_{i(jk)}-C^{2}_{ik}},\;i\neq j\neq k\;\text{and}\;i,j,k=1,2,3$ (10)
Partial tangles help classify the entanglement structure in a three-qubit state datta2018 :
  • (i)
    If all partial tangles vanish, the state is fully separable.
  • (ii)
    If at least one partial tangle is zero, the state is biseparable.
  • (iii)
    If all are non-zero, the state exhibits genuine tripartite entanglement.
If all the partial tangles are zero, then the three qubit state is separable. If atleast one partial tangle is zero, then the state is biseparable. If each partial tangle is non-zero, then the state is genuine entangled datta2018 .
II.4 Concurrence fill
Xie and Eberly introduced a triangle measure of entanglement for three-qubit systems, named as, concurrence fill eberly2021 . It has the advantage of being nonincreasing under local quantum operations and classical communications (LQCC). This measure is based on the entanglement polygon inequality which states that one entanglement cannot increase the sum of the other two qian2018 ; zhu2015 , i.e.,
$C^{2}_{i|jk}\leq C^{2}_{j|ki}+C^{2}_{k|ij}$ (11)
where a subscript $i$ refers to the system’s $i$th qubit. A concurrence triangle is the geometrical interpretation of entanglement polygon inequality eberly2021 .
[…]
Figure 1: The above figure shows the concurrence triangle for the three-qubits A, B, and C. The edges of the triangle represents the squared concurrences between bipartitions.
Concurrence fill, denoted by $C_{F}$, is a genuine tripartite entanglement measure that represents the area of the concurrence triangle. For a three qubit state, $\rho_{ABC}$, concurrence fill is given as
$C_{F}(\rho_{ABC})=\left[\frac{16}{3}Q(Q-C^{2}_{A|BC})(Q-C^{2}_{B|AC})\right.\\ \left.\times(Q-C^{2}_{C|AB})\right]^{1/4},\\ \text{where}\quad Q=\frac{1}{2}\left(C^{2}_{A|BC}+C^{2}_{B|AC}+C^{2}_{C|AB}\right)$ (12)
$C_{F}$ is zero for all separable and biseparable states and is non-zero for genuine entangled states (GHZ and W states). It is given as
$C_{F}(|GHZ\rangle)=1\;\text{and}\;C_{F}(|W\rangle)=\frac{8}{9}$ (13)
$C_{F}(|GHZ\rangle)=1$ and Also, $C_{F}$ is monotonic under LQCC.
Concurrence fill can be generalized for mixed states via convex roof extension. It is given as
$C_{F}(\rho_{ABC})=min_{\{p_{i},|\psi_{i}\rangle\}}\sum_{i}{p_{i}C_{F}(|\psi_{i}\rangle})$ (14)
where the minimum is taken over all convex decompositions of $\rho_{ABC}$.
III Classification of GHZ and W-class of states
In this section, we delve into the geometric entanglement measure known as concurrence fill for three-qubit pure states. We begin by expressing concurrence fill in terms of the tangle and partial tangle, demonstrating that this formulation remains invariant under different basis choices. Subsequently, we introduce a method to classify three-qubit states into the GHZ and W classes using concurrence fill. To lay the groundwork, let us first examine the basis representation of three-qubit states.
It is known that any three-qubit pure state can be written in canonical form, where the positive Schmidt coefficients contain all the information about the non-local properties of quantum states acin2001 . Canonical forms of three-qubit pure states are classified based on minimal number of local basis product states, i.e., minimal number of non-local parameters, needed for the specification of a state. In three qubit system, there are three inequivalent classes of local basis product states given as follows.
$\{|000\rangle,|001\rangle,|010\rangle,|100\rangle,|111\rangle\}$ (15)
$\{|000\rangle,|001\rangle,|100\rangle,|110\rangle,|111\rangle\}$
$\{|000\rangle,|001\rangle,|011\rangle,|100\rangle,|111\rangle\}$ (16)
$\{|000\rangle,|010\rangle,|100\rangle,|101\rangle,|111\rangle\}$
$\{|000\rangle,|100\rangle,|101\rangle,|110\rangle,|111\rangle\}$
$\{|000\rangle,|010\rangle,|011\rangle,|110\rangle,|111\rangle\}$ (17)
$\{|000\rangle,|001\rangle,|011\rangle,|101\rangle,|111\rangle\}$
where the three sets of states in (16) and (17) are related by permutation.
Acin et. al have shown that any three-qubit pure state $|\psi\rangle$ can be expressed as acinprl2001
$|\psi_{ABC}\rangle=\lambda_{0}|000\rangle+\lambda_{1}e^{i\theta}|100\rangle+\lambda_{2}|101\rangle$
$+\lambda_{3}|110\rangle+\lambda_{4}|111\rangle$ (18)
where $\lambda_{i}\geq 0$, $\sum_{i=0}^{4}\lambda_{i}=1$ and $\theta\in[0,\pi]$.
Note that the pure state $|\psi\rangle$ in (18), belongs to the class of symmetric states in (17).
Now we express the concurrence fill of a three-qubit state in terms of three-tangle and partial tangle. Let us first consider the state $|\psi_{ABC}\rangle$ given in (18).
The concurrences and tangle between any two qubits are given by
$\tau_{\psi}$ $=$ $4\lambda_{0}^{2}\lambda_{4}^{2}$ (19)
$C^{2}_{AC}$ $=$ $4\lambda_{0}^{2}\lambda_{2}^{2}$ (20)
$C^{2}_{AB}$ $=$ $4\lambda_{0}^{2}\lambda_{3}^{2}$ (21)
$C^{2}_{BC}$ $=$ $4(\lambda_{2}^{2}\lambda_{3}^{2}-\lambda_{1}^{2}\lambda_{4}^{2})^{2}$ (22)
The partial tangles for the state $|\psi_{ABC}\rangle$ are given as
$\tau^{2}_{ij}=C^{2}_{ij}+\tau_{\psi}\;\;\;\forall\;\;i,j=1,2,3$ (23)
After evaluating the relevant expressions, the concurrence fill of the state $|\psi_{ABC}\rangle$ can be expressed as
$C_{F}(|\psi_{ABC}\rangle)=\left(\frac{16}{3}\prod_{i<j}(\tau_{ij}^{2}-\frac{\tau_{\psi}}{2})\sum_{i<j}(\tau_{ij}^{2}-\frac{3\tau_{\psi}}{2})\right)^{1/4}$ (24)
for $i,j=1,2,3$
It is interesting to note that the above expression of the concurrence fill is independent of the choice of basis, i.e., for any basis given in (15), (16) and (17), area under the concurrence triangle remains same.
Now we are in a position to state a few results based on classification of GHZ and W class of states.
Theorem. The following inequality holds for W-class of states.
${[C_{F}(|\psi_{ABC}\rangle)]^{4}}\geq k\;\tau^{2}_{BC}$ (25)
where $k=(\tau^{2}_{AB}+C^{2}_{AB})^{4/3}(\tau^{2}_{AC}+C^{2}_{AC})^{4/3}(\tau^{2}_{BC}+C^{2}_{BC})^{1/3}$
Proof.
Using (24) we have,
$\begin{split}C_{F}(|\psi_{ABC}\rangle)=\bigg{(}\frac{1}{3}(\tau_{AB}^{2}+C^{2}_{AB})(\tau_{AC}^{2}+C^{2}_{AC})(2\tau_{BC}^{2}-\tau_{\psi})\\ (2C^{2}_{BC}+C^{2}_{AC}+C^{2}_{AB}+\tau^{2}_{AB}+\tau^{2}_{AC}+\tau_{\psi})\bigg{)}^{1/4}\end{split}$ (26)
$\begin{split}C_{F}(|\psi_{ABC}\rangle)\geq\bigg{(}\frac{1}{3}(\tau_{AB}^{2}+C^{2}_{AB})(\tau_{AC}^{2}+C^{2}_{AC})\\ (\tau_{BC}^{2}-\tau_{\psi})\sum_{i<j}(\tau_{ij}^{2}+C^{2}_{ij})\bigg{)}^{1/4}\end{split}$ (27)
Using AM-GM inequality on above equation, we have
$\begin{split}C_{F}(|\psi_{ABC}\rangle)\geq((\tau_{AB}^{2}+C^{2}_{AB})(\tau_{AC}^{2}+C^{2}_{AC})\\ (\tau_{BC}^{2}-\tau_{\psi})\prod_{i<j}(\tau^{2}_{ij}+C_{ij}^{2})^{1/3})^{1/4}\end{split}$ (28)
i.e.,
$\begin{split}C_{F}(|\psi_{ABC}\rangle)\geq(\tau_{AB}^{2}+C^{2}_{AB})^{1/3}(\tau_{AC}^{2}+C^{2}_{AC})^{1/3}\\ (\tau^{2}_{BC}+C_{BC}^{2})^{1/12}(\tau_{BC}^{2}-\tau_{\psi})^{1/4}\end{split}$ (29)
Proof follows from the inequality (29). ∎
Corollary. Let $|\psi_{ABC}\rangle$ be a three-qubit genuinely entangled state. If the inequality in (25) is violated, i.e.,
${[C_{F}(|\psi_{ABC}\rangle)]^{4}}<k\;\tau^{2}_{BC}$ (30)
then, $|\psi_{ABC}\rangle$ belongs to GHZ class of states.
Example 1.
Consider the three-qubit state $|\phi\rangle=c|001\rangle+d|010\rangle+f|100\rangle+e|011\rangle$, where $|c|^{2}+|d|^{2}+|f|^{2}+|e|^{2}=1$
  • Case 1.
    For $c=0$, $e=\frac{1}{2}$, $d,f\neq 0$ we have
    $|\phi_{1}\rangle=d|010\rangle+f|100\rangle+\frac{1}{\sqrt{5}}|011\rangle$. Note that $\tau(|\phi_{1}\rangle)=e^{2}f^{2}\neq 0$ which implies that $|\phi_{1}\rangle$ is GHZ state. Now using the criteria given in above result, we see that the inequality (30) is violated which implies that $|\phi_{1}\rangle$ belongs to GHZ class.
    Figure 2: The blue and yellow lines represent the lhs and rhs of the inequality (30) for the states $|\phi_{1}\rangle$.
  • Case 2.
    For $e=0$, $c=\frac{1}{\sqrt{5}}$, $d,f\neq 0$, $|\phi_{2}\rangle=c|001\rangle+d|010\rangle+f|100\rangle$. In this case, we have
    $[C_{F}(|\phi_{2}\rangle)]^{4}\geq k\;\tau^{2}_{BC}$ (31)
    Thus, using (25), $|\phi_{2}\rangle$ belongs to W class. Also, since the three-tangle $\tau(|\phi_{2}\rangle)=e^{2}f^{2}=0$, and three partial tangles are non-zero, the state $|\phi_{2}\rangle$ belongs to W class of states.
    Figure 3: The blue and yellow lines represent the lhs and rhs of the inequality (31) for the states $|\phi_{2}\rangle$.
IV Concurrence Fill for a Rank-two Mixture of states
In this section, we study the rank-2 mixture of generalized GHZ and W states. Firstly we derive the minimum value of the concurrence fill for any pure state superposition of generalized GHZ and W states. Next, we formulate the concurrence fill for the mixed states using optimal decomposition.
Let us consider the following family of mixed three-qubit states.
$\rho(p)=p|gGHZ\rangle\langle gGHZ|+(1-p)|gW\rangle\langle gW|$ (32)
where $0\leq p\leq 1$; $|gGHZ\rangle$ and $|gW\rangle$ are the generalized GHZ and W states, respectively, given as follows.
$|gGHZ\rangle$ $=$ $a|000\rangle+b|111\rangle$ (33)
$|gW\rangle$ $=$ $c|001\rangle+d|010\rangle+f|100\rangle$ (34)
where $|a|^{2}+|b|^{2}=1$ and $|c|^{2}+|d|^{2}+|f|^{2}=1$.
IV.1 Concurrence fill for eigenstates of $\rho(p)$
In order to find the concurrence fill for mixed states, let us first study the pure states that are superpositions of the eigenstates of $\rho$. Note that these states belong to the basis given in (15).
$|\psi_{p,\phi}\rangle=\sqrt{p}|gGHZ\rangle-\sqrt{1-p}\;e^{i\phi}|gW\rangle$ (35)
Three-tangle for the pure state $|\psi_{p,\phi}\rangle$ is given as follows oster2006 ; oster2008 ; njp2008 .
$\tau(|\psi_{p,\phi}\rangle)=4|p^{2}a^{2}b^{2}-4\sqrt{p(1-p)^{3}}e^{3i\phi}bcdf|$ (36)
Using (12), we derive the concurrence fill of the state $|\psi_{p,\phi}\rangle$ and it is given as follows:
$C_{F}(|\psi_{p,\phi}\rangle)=\left(\frac{256}{3}|xyz(x+y+z)|\right)^{1/4}$ (37)
where
$x$ $=$ $p^{2}a^{2}b^{2}+2p(1-p)e^{2i\phi}f^{2}b^{2}+2e^{4i\phi}(1-p)^{2}c^{2}d^{2}$
$y$ $=$ $p^{2}a^{2}b^{2}+2p(1-p)e^{2i\phi}b^{2}d^{2}+2e^{4i\phi}(1-p)^{2}e^{2}f^{2}$
$z$ $=$ $p^{2}a^{2}b^{2}+2p(1-p)e^{2i\phi}b^{2}c^{2}+2e^{4i\phi}(1-p)^{2}d^{2}f^{2}$
We note that
$C_{F}(|gGHZ\rangle)$ $=$ $4|a^{2}b^{2}|$ (39)
$C_{F}(|gW\rangle)$ $=$ $8|cdf|\left(\frac{|c^{2}d^{2}+d^{2}f^{2}+c^{2}f^{2}|}{3}\right)^{1/4}$ (40)
Now we derive a lower bound of the concurrence fill for the pure eigenstates of the form (35), which is appended in the following result.
Result. For the three-qubit pure state $|\psi_{p,\phi}\rangle$ given in (35), the following inequality holds.
$C_{F}(|\psi_{p,\phi}\rangle)\geq C_{F}^{(min)}(|\psi_{p_{min},\phi}\rangle)$ (41)
where $p_{min}=\frac{1}{1+3b^{2}}$, and
[…]
represents the lower bound of the concurrence fill for the state $|\psi_{p,\phi}\rangle$.
Proof.
In order to derive the lower bound of the concurrence fill, first we choose $\phi=2n\pi$ such that $e^{i\phi}=1$ where $n=1,2,3,...$. Now our aim is to find the value of $p$ for which the state $|\psi_{p,\phi}\rangle$ attains minima. From (37), we have
$\frac{dC_{F}(|\psi_{p,\phi}\rangle)}{dp}=\frac{x^{\prime}}{x}+\frac{y^{\prime}}{y}+\frac{z^{\prime}}{z}+\frac{(x^{\prime}+y^{\prime}+z^{\prime})}{x+y+z}$ (43)
where $x^{\prime}=\frac{dx}{dp},\;y^{\prime}=\frac{dy}{dp},\;z^{\prime}=\frac{dz}{dp}$ and $x,y,z$ are defined in (LABEL:xyz).
After some calculations, we find that $\frac{dC_{F}(|\psi_{p,\phi}\rangle)}{dp}=0$, when
$p=p_{1}=p_{2}=p_{3}$ (44)
where
$p_{1}$ $=$ $\frac{2c^{2}d^{2}-f^{2}b^{2}}{a^{2}b^{2}+2(c^{2}d^{2}-f^{2}b^{2})}$ (45)
$p_{2}$ $=$ $\frac{2c^{2}f^{2}-b^{2}d^{2}}{a^{2}b^{2}+2(c^{2}f^{2}-b^{2}d^{2})}$ (46)
$p_{3}$ $=$ $\frac{2d^{2}f^{2}-b^{2}c^{2}}{a^{2}b^{2}+2(d^{2}f^{2}-b^{2}c^{2})}$ (47)
Now, we find that (44) holds when $|a|^{2}+|b|^{2}=1$ and $c=d=f=\frac{1}{\sqrt{3}}$. Hence, $p_{min}=\frac{1}{1+3b^{2}}$ is the point of minima with $10.25<p_{min}<$ 1, at which the concurrence fill is given by
$C_{F}^{(min)}(|\psi_{p_{min},\phi}\rangle)=\frac{4b^{2}}{1+3b^{2}}\;\text{where}\;0<b<1$ (48)
Hence, we obtain a lower bound of the concurrence fill for the pure states of the form (35) given as
$C_{F}(|\psi_{p,\phi}\rangle)\geq\frac{4b^{2}}{1+3b^{2}}\;\text{where}\;0<b<1$ (49)
Example 2.
Consider the symmetric GHZ state and symmetric W state in (35). In this case, we have $a=b=\frac{1}{\sqrt{2}}$ and $c=d=f=\frac{1}{\sqrt{3}}$ which gives $p_{min}=\frac{2}{5}$, and the concurrence fill is given as $C_{F}(|\psi_{\frac{2}{5},\phi}\rangle)=\frac{4}{5}$. Hence, the lower bound obtained in (49) conincides with the actual expression of concurrence fill for symmetric mixture of GHZ and W class of states.
IV.2 Concurrence fill for mixed states
The convex roof construction is a widely used method for quantifying the entanglement of mixed states. The ensemble that minimizes the average value of an entanglement measure under this construction is referred to as the optimal ensemble. While the concurrence of two-qubit mixed states admits an analytical solution, in general, determining the convex roof is highly nontrivial. Nonetheless, several effective approaches have been developed to estimate it, including techniques that provide upper and lower bounds as useful approximations oster2008 ; guhne2015 .
Let us consider the mixed state
$\rho=\sum_{i}p_{i}\Omega_{i},\;\Omega_{i}=\frac{|\phi_{i}\rangle\langle\phi_{i}|}{\langle\phi_{i}|\phi_{i}\rangle}$ (50)
The concurrence fill of the mixed state $\rho$ can be generalized as convex roof extension of the concurrence fill of pure states,
$C_{F}(\rho)=min_{\{p_{i},\Omega_{i}\}}\sum_{i}p_{i}C_{F}(\Omega_{i})$ (51)
Our aim is to estimate the concurrence fill of the generalized mixture of GHZ and W class of states. Let us consider the mixed state of the form $\rho(p)$ given in (32). In njp2008 , the optimal decomposition of the zero tangle mixed states $\rho(p)$ has been derived. It is given as follows.
[…]
where
$\hat{\rho}(p)=\frac{1}{3}\sum_{k=0}^{2}|\psi_{p,\frac{2k\pi}{3}}\rangle\langle\psi_{p,\frac{2k\pi}{3}}|$ (53)
and
$p_{0}=\frac{(16c^{2}d^{2}f^{2})^{1/3}}{(a^{4}b^{2})^{1/3}+(16c^{2}d^{2}f^{2})^{1/3}}$ (54)
with $|\psi_{p,\phi}\rangle$ is defined in (35) and $\Omega_{i}$ is defined in (50).
We have $\tau(\rho(p))=0$ when $0\leq p\leq p_{0}$. The concurrence fill of the state $\hat{\rho}(p)$ is given as
$C_{F}(\hat{\rho}(p))\leq\frac{1}{3}\left[C_{F}(|\psi_{p,0}\rangle)+C_{F}(|\psi_{p,\frac{2\pi}{3}}\rangle)+C_{F}(|\psi_{p,\frac{4\pi}{3}}\rangle)\right]$ (55)
Hence, for any mixed state $\rho(p)$ with $\tau(\rho(p))=0$ where $0\leq p\leq p_{0}$, the upper bound of the concurrence fill is given as follows.
$C_{F}(\rho(p))\leq C^{ub}_{F}(\rho(p))$ (56)
where
$C^{ub}_{F}(\rho(p))=\frac{p}{p_{0}}C_{F}(\hat{\rho}(p))+\frac{p_{0}-p}{p_{0}}C_{F}(|gW\rangle)$ (57)
with $C_{F}(\hat{\rho}(p))$ is derived in (55) and $C_{F}(|gW\rangle)$ has been derived in (40).
Example 3.
[…]
Figure 4: The blue curve represents the upper bound of the concurrence fill given in (57) for the zero-tangle states in the generalized mixture $\rho(p)$ with $0\leq p\leq p_{0}$.
V Conclusion
In this work, we addressed the challenge of detecting and classifying three qubit genuine entangled states, which remains elusive under many conventional measures despite experimental realizations. To achieve the task, we study the concurrence fill for pure as well as mixed three-qubit states. We began by reformulating the concurrence fill in terms of three tangle and partial tangle. This characterization is important because partial tangle can be measured using expectation value of measurement operators which makes them operationally meaningful. We further employ our formulation to classify GHZ and W class of states. Since convex roof extension is a computationally challenging task, we applied our framework to analyze a rank-2 mixture of generalized GHZ and W-class states, which serves as a representative for nontrivial quantum mixtures. For the eigenstates of the resulting generalized mixed states, we derived an explicit expression for concurrence fill and established an upper bound for the zero tangle mixed states.
These findings not only enhance the utility of concurrence fill as a diagnostic tool for detecting genuine three qubit entangled states, but also open the door to future studies on geometric characterizations of quantum measures in multipartite systems.
VI Acknowledgement
The author acknowledges initial academic guidance provided by Dr. Satyabrata Adhikari during the early stages of this work.
VII DATA AVAILABILITY STATEMENT
Data sharing not applicable to this article as no datasets were generated or analysed during the current study.